91 169 305 330 335 389 393 394 402 410 420 449 452 455 459 465...

91
169
305
330
335
389
393
394
402
410
420
449
452
455
459
465
468
471
472
474
475
479
481
486
487
505
508
509
511
512
517
519
530
537
For this data set is 44.6 the correct class limit to use even if the upper limit ends in 537.9?
Answer #1

Difference = 537-91= 446

if you want 10 class differences then the width of the class interval = 446/10= 44.6 next integer = 45

which is not the class limit

The lower class limit will be 91 and the upper class will be 91+45 = 136.( 91 -- 136) first class interval and its class limit

So, 44.6 is not a correct class limit.